# MAMBO-G: Magnitude-Aware Mitigation for Boosted Guidance

### v3

@@ 1 Introduction
Generative models have made significant progress in creating images and videos from text (Song et al., 2021; Dhariwal & Nichol, 2021; Peebles & Xie, 2023; Ho et al., 2022; Esser et al., 2023). A key technique they use is classifier-free guidance (CFG) (Ho & Salimans, 2021), which adjusts the model’s output to better match the text prompt. However, using strong guidance can sometimes reduce stability, potentially leading to issues like oversaturated colors or unnatural structures (Karczewski et al., a; Sadat et al., 2024). These problems are often more noticeable in modern, high-dimensional models, where using guidance scale without careful adjustment may strongly affect visual quality.
@@ 2.1 Guidance Strategies in Diffusion Models
Various methods aim to improve guidance in diffusion models. CFG++ (Chung et al., ) treats guidance as a manifold-constrained inverse problem (Karczewski et al., 2025), while CFG Schedulers (Xi et al., ) and Apply Guidance in Interval (Kynkäänniemi et al., 2024) optimize time-dependent strength; similarly, Stage-Wise Dynamics (Jin et al., 2025) investigates varying guidance requirements across different generation stages. ReCFG (Xia et al., 2025) and TFG (Ye et al., 2024) focus on correcting expectation shifts, a goal shared by Rectified CFG++ (Saini et al., 2025) and CFG-EC (Yang et al., 2025) which propose mechanisms to rectify guidance errors and improve consistency. Recent approaches explore dynamic adaptation: $\mathbf{S^{2}}$-Guidance (Chen et al., 2025) uses stochastic block-dropping, REG (Gao et al., 2025) optimizes scaled joint distributions, FBG (Koulischer et al., 2025) uses feedback on conditional informativeness, and Foresight Guidance (Wang et al., 2025) frames CFG as a fixed-point iteration with short inner loops. Complementary to these, Prompt-Aware Guidance (Zhang & Li, 2025) adapts strength based on prompt complexity, Learning-to-Guide (Galashov et al., 2025) employs meta-learning for optimal strategies, and Saddle-Free Guidance (Yeats et al., 2025) navigates the optimization landscape to avoid saddle points. Density Guidance (Karczewski et al., b) extends flow matching by incorporating explicit log-density control to steer generation trajectories. While it offers a rigorous theoretical unification of prior heuristics via Score Alignment, the method is computationally expensive. Specifically, estimating the divergence requires Jacobian-Vector Products (JVP), which doubles the inference cost and introduces estimation variance. In Flow Matching models (Esser et al., ; Lipman et al., 2023; Fan et al., 2025a; Liu et al., 2023; Gao et al., 2024), CFG-Zero* (Fan et al., 2025b) compensates for velocity errors.

### v4

@@ 1 Introduction
Generative models have made significant progress in creating images and videos from text (Song et al., 2021; Dhariwal and Nichol, 2021; Peebles and Xie, 2023; Ho et al., 2022; Esser et al., 2023). A key technique they use is classifier-free guidance (CFG) (Ho and Salimans, 2021), which adjusts the model’s output to better match the text prompt. However, using strong guidance can sometimes reduce stability, potentially leading to issues like oversaturated colors or unnatural structures (Karczewski et al., ; Sadat et al., 2024). These problems are often more noticeable in modern, high-dimensional models, where using guidance scale without careful adjustment may strongly affect visual quality.
@@ 2.1 Guidance Strategies in Diffusion Models
Various methods aim to improve guidance in diffusion models. CFG++ (Chung et al., ) treats guidance as a manifold-constrained inverse problem (Karczewski et al., 2025), while CFG Schedulers (Xi et al., ) and Apply Guidance in Interval (Kynkäänniemi et al., 2024) optimize time-dependent strength; similarly, Stage-Wise Dynamics (Jin et al., 2025) investigates varying guidance requirements across different generation stages. ReCFG (Xia et al., 2025) and TFG (Ye et al., 2024) focus on correcting expectation shifts, a goal shared by Rectified CFG++ (Saini et al., 2025) and CFG-EC (Yang et al., 2025) which propose mechanisms to rectify guidance errors and improve consistency. Recent approaches explore dynamic adaptation: $\mathbf{S^{2}}$-Guidance (Chen et al., 2025) uses stochastic block-dropping, REG (Gao et al., 2025) optimizes scaled joint distributions, FBG (Koulischer et al., 2025) uses feedback on conditional informativeness, and Foresight Guidance (Wang et al., 2025) frames CFG as a fixed-point iteration with short inner loops. Complementary to these, Prompt-Aware Guidance (Zhang and Li, 2025) adapts strength based on prompt complexity, Learning-to-Guide (Galashov et al., 2025) employs meta-learning for optimal strategies, and Saddle-Free Guidance (Yeats et al., 2025) navigates the optimization landscape to avoid saddle points. Density Guidance (Karczewski et al., ) extends flow matching by incorporating explicit log-density control to steer generation trajectories. While it offers a rigorous theoretical unification of prior heuristics via Score Alignment, the method is computationally expensive. Specifically, estimating the divergence requires Jacobian-Vector Products (JVP), which doubles the inference cost and introduces estimation variance. In Flow Matching models (Esser et al., ; Lipman et al., 2023; Fan et al., 2025a; Liu et al., 2023; Gao et al., 2024), CFG-Zero* (Fan et al., 2025b) compensates for velocity errors.